40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1235     Entry time: Fri Jan 16 18:33:54 2009
Author: Yoichi 
Type: Summary 
Category: Computers 
Subject: c1lsc rebooted to fix 16Hz glitches 
Kakeru, Yoichi

There were 16Hz harmonics in the PD3 and PD4 channels even when there is no light falling on it.
Actually, even when the connection to the ADC was removed, the 16Hz noise was still there.

Rob suggested that this might be digital problem, because data is sent to the daq computer very 1/16 of a second.

We restarted c1lsc and the problem went away.
ELOG V3.1.3-